for what value of k does the equation kt^2 + 5t +2 = 0, have exactly one real solution for t?
Answer by stanbon(75887)

for what value of k does the equation kt^2 + 5t +2 = 0, have exactly one real solution for t?
----
Quadratic with a = k ; b = 5 and c = 2
---
Discriminant:
b^2 -4ac = 5^2 -4k*2
---
Solve:
5^2-4k*2 = 0
8k = 25
k = 25/8
------------------
Cheers,
Stan H.